OpenStudy (anonymous):

Indicate the equation of the line through (2, -4) and having slope of 3/5.

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (zale101):

Slope formula (plug m 3/5) and plug (2,-3) for x,y Y=mx+b Point slope formula y-y1=m(x-x2) Plug to this one Which formula you prefer?

OpenStudy (alexandervonhumboldt2):

Y=mx+b is bettter

OpenStudy (anonymous):

 this is how it has to be